OpenStudy (anonymous):

I travel to 900km on one particular day, on my return trip my speed is such that it takes me 2hrs less to get back. How long are the individual trips and whats my return speed?

OpenStudy (anonymous):

\[ v_1 = \frac{900km}{t_1} \] \[v_2 = \frac{900km}{t_1-2}\] focus on the \(v_1\) \[t_1 = \frac{900km}{v_1}\] put this equation into \(v_2\) \[ v_2 = \frac{900km}{\frac{900km}{v_1}-2}\] This is how far I got... do you have some more information?

OpenStudy (amistre64):

distance = distance right?

OpenStudy (amistre64):

lol ... making up questions tends to be harder than it looks ;)

OpenStudy (amistre64):

you gain a certain appreciation for the questions in a textbook after doing it for awhile; its a conspiracy i tell ya :)

OpenStudy (anonymous):

Lol no actuali I should have said V2 is 20km/h hire. ignore my last comment

OpenStudy (anonymous):

yeah, has a habit of actually having an answer :P

OpenStudy (anonymous):

I remember this from high school, about 4yrs ago

OpenStudy (anonymous):

90